25.00 mL of 6.00 M H2SO4 solution is diluted with water to a volume of 2.000L. What is the [H+] and [OH-] of the resulting solution?


1
Expert's answer
2021-12-14T03:43:04-0500

Solution:

Convert mL to L:

(25.00 mL) × (1 L / 1000 mL) = 0.025 L


Calculate the concentration of the diluted H2SO4 solution:

(6.00 M H2SO4) × (0.025 L / 2.000 L) = 0.075 M

CM(H2SO4) = 0.075 M


H2SO4 is a strong acid which dissociates as follows:

H2SO4 → 2H+ + SO42-

Therefore,

[H+] = 2 × CM(H2SO4) = 2 × (0.075 M) = 0.15 M

[H+] = 0.15 M


Kw = [H+] × [OH-] = 1.0×10-14

Therefore,

[OH-] = Kw / [H+] = (1.0×10-14) / (0.15) = 6.67×10-14

[OH-] = 6.67×10-14 M


Answer:

[H+] = 0.15 M

[OH-] = 6.67×10-14 M